# Quantum Neural Networks in Practice: A Comparative Study with Classical Models from Standard Data Sets to Industrial Images

### v1

@@ Abstract
Image classification tasks are among the most prominent examples that can be reliably solved by classical machine learning models. In this study, we compare the performance of randomized classical and quantum neural networks as well as classical and quantum-classical hybrid convolutional neural networks for the task of binary image classification. To this end, we employ various data sets of increasing complexity — (i) an artificial hypercube dataset, (ii) MNIST handwritten digits, and (iii) real-world industrial images from laser cutting machines. We analyze the performance of the employed quantum models with respect to correlations between classification accuracy and various hyperparameters. For the random quantum neural networks, we additionally compare their performance with some known literature models and how top-performing models from one data set perform on the others. In general, we observe fairly similar performances of classical and quantum or hybrid models. Our study provides an industry perspective on the prospects of quantum machine learning for practical image classification tasks.
@@ I Introduction
Quantum machine learning [1, 2] is one of the various subfields of quantum computing promising advantages over its classical counterpart, for instance due to proposals regarding more efficient data encoding [3, 4], improved model expressivity [5, 6, 7], prediction accuracy [8, 9, 10] or by saving computational resources [11, 12]. These theoretical proposals have been backed up by claims of experimental quantum advantage in learning tasks [13, 14]. Despite undeniable progress in the field of quantum machine learning — both on the algorithmic as well as the hardware side — doubts about the field’s focus on quantum advantage have been previously raised [15]. Ultimately, this culminated in a recent work questioning the significance of quantum-machine-learning benchmark studies [16], given their predominant use of overly simple data sets paired with insufficiently-performing quantum hardware. Other works [17, 18, 19] even suggest that much of what was commonly believed to be classically hard or intractable to simulate might be efficiently simulable on classical hardware after all. While this may require the field to reevaluate what “quantum” means for quantum machine learning and how a future quantum computer may use “quantumness” advantageously for learning tasks, both Refs. [16, 17] independently call for more rigorous benchmarks employing less trivial data sets in order to assess the potential of quantum machine learning.
@@
In this work, we benchmark the performance of classical and quantum machine learning models in classifying three data sets of increasing data complexity. The data sets include (i) an artificial data set based on linearly splitting a (hyper-)cube in two, (ii) MNIST’s handwritten digits [20] as an easy but well-studied data set and (iii) images from laser cutting machines as a challenging data set from the industry. Note that we use the first two data sets merely as a means to benchmark and validate our employed methodologies, since — as mentioned in Refs. [16, 18] — the potential insights about the power and usefulness of quantum machine learning that can be drawn from simple data sets are very limited. However, we still believe them to be useful for validating the results we obtain for the practically-relevant industrial data set on which we put the main focus when drawing conclusions on quantum machine learning from an industry perspective. Moreover, since we are also interested in testing and assessing the cross-data-set performance of some quantum models, we deem it beneficial to have results for various data sets with different levels of complexity available.
@@
In terms of methodologies, we employ both classical and quantum neural networks [5] as well as classical and quantum-classical hybrid convolutional neural networks [21] to classify any of the mentioned data sets. Our decision for parametrized quantum circuits is motivated by their high expressivity [5] which is amplified by techniques like data re-uploading [22, 23] and trainable data embeddings [24] while at the same time keeping quantum hardware requirements moderate. This is in contrast to some early suggestions for quantum machine learning models based on amplitude encoding [4, 3, 25], which, despite being qubit-efficient, were later deemed too resource-demanding to run on any near-term device due to their excessive need of quantum gates for accurate state preparation [6]. In contrast to these advantages, parametrized quantum circuits suffer from barren plateaus [26], which prevents their efficient training at scale, although much work has been put into mitigating or avoiding them, see e.g. Refs. [27, 28, 29, 30] to name but a few. Some quantum models, such as quantum convolutional neural networks [31], provably avoid barren plateaus [32]. Unfortunately, the latter are also rather resource-demanding to qualify for near-term applicability and practical simulability within this study, which is why we resort to quantum-classical hybrid convolutional neural networks [21] instead. Conceptually, parametrized quantum circuits as well as quantum-classical hybrid convolutional neural networks belong to the realm of variational quantum algorithms [33], where the execution of quantum circuits is carried out on actual quantum hardware, while the calculation of iteratively-updated circuit parameters is performed on classical computers. Due to these choices, our study reflects a current perspective on what quantum machine learning methods, suitable for present-day or near-term quantum devices, offer for classifying industry-level image data.
@@ II Data sets
Figure 1: Overview of data sets with $N=500$ and dimension $d=2$ in panels (a)-(e), and the corresponding learned functions by classical and quantum neural networks in panels (f)-(j) and panels (k)-(o), respectively. The small neural networks in between panels (a)-(e) and (f)-(j) show the best performing architectures (from among the set of randomly generated neural networks). The small quantum neural networks in between panels (a)-(e) and (k)-(o) show the respective best performing quantum architectures (from among the set of randomly generated quantum neural networks). Data points represented by circles (stars) in panels (a)-(o) correspond to the training (validation) data set. In addition, panels (p) and (q) show the architecture for the convolutional autoencoder (CAE) used to reduce the data dimensionality for the images of the second and third data set, respectively, as well as a representative illustration of the original and reconstructed images depending on the latent dimension $d$.
@@ II.1 Data set 1 — linear split
The first data set is an artificial one and the only one not based on images. It is chosen with the premise of simplicity in mind and such that the ideal function, which classifies the data set perfectly, is known in advance. To this end, we choose to consider a $d$-dimensional hypercube, defined by the set $[0,1]^{d}\subseteq\mathbb{R}^{d}$, and linearly split it by the plane orthogonal to the normal vector $\boldsymbol{n}=(1,1,\dots,1)^{\top}/\sqrt{d}\in\mathbb{R}^{d}$ which passes through the hypercube’s center $(1,1,\dots,1)^{\top}/2\in\mathbb{R}^{d}$. In order to generate training and validation data sets, we first place $N$ data points randomly within the hypercube such that exactly $N/2$ points are located within each of the two sectors separated by the hyperplane. These two sectors define the two labels that classify each data point. Afterwards, we split the $N$ data points into training (validation) data sets consisting of $N_{\mathrm{train}}=0.8N$ ($N_{\mathrm{test}}=0.2N$) of the total data points. We perform this splitting such that the number of data points with each label within each subset is identical. In order to study the effect of the hypercube’s dimension $d$ as well as the data set’s size $N$, we generate versions of the hypercube data set with $d\in\{2,4,8,16,32,64\}$ and $N\in\{200,500,1000,2000\}$. An example with $d=2$ and $N=500$ is shown in Fig. 1 (c).
This hypercube data set has recently also been studied in a related work exploring the performance of quantum machine learning methods [16]. In this work, the hypercube data turned out to be rather difficult to classify by means of quantum machine learning methods despite being a trivial problem for classical machine learning. Hence, we use it in our study as an ideal means to validate and benchmark our methodologies, introduced in Sec. III, and to check whether it can reproduce those results qualitatively.
@@ II.2 Data set 2 — handwritten digits
As a second data set, we take the well-studied handwritten digits of the MNIST database [20] as another means to test our methodologies. However, in order to keep the classification problem binary, we restrict our study to images of 00’s and $1$’s. Similar to the hypercube’s data sets described in Sec. II.1, we generate different versions of the MNIST data set with $N\in\{200,500,1000,2000\}$ total images taken randomly from the $60000$ available images. We again ensure that each generated data set has an equal number of 00’s and $1$’s and that this ratio is maintained once the total data set is split into training and validation sets with $N_{\mathrm{train}}=0.8N$ and $N_{\mathrm{test}}=0.2N$ data points, respectively.
@@ II.3 Data set 3 — industrial images
In contrast to Fig. 2, the slat images we use in the following are already preprocessed and cut such as to focus on a single slat socket which may or may not contain a slat. The classification problem is thus binary with two difficulties arising due to varying camera angles from which the slat sockets are being photographed as well as the different condition of the slats, cf. Fig. 2. To ease data generation, TRUMPF has developed a model that generates synthetic slat socket images given the CAD model of the entire laser cutting machine and textures derived from real-world images using the Blender rendering engine. This way, manual labeling effort of real-world images is unnecessary and labeling errors are ruled out. Moreover, it allows to generate data sets with large image numbers. Identically to the other data sets, we consider different versions of this industry data set with $N\in\{200,500,1000,2000\}$ total images, which are afterwards split into training and validation data sets with $N_{\mathrm{train}}=0.8N$ and $N_{\mathrm{test}}=0.2N$ data points, respectively. For these two sets, we ensure again the same class distribution as the original data set, as well as similar number of images of slat sockets photographed with different camera angles and conditions. All images are gray-scale and normalized to $200\times 80$ pixels independent of the photo angle.
@@ III.1.1 Classical neural networks
In order to generate random classical neural network architectures that allow processing of data vectors $\boldsymbol{x}_{i}\in\mathbb{R}^{d}$, there are various hyperparameters that can be randomly determined. Among the most prevalent ones are the number of dense, i.e., fully-connected, layers and the number of neurons per layer as well as the employed activation functions. We decide to randomize the first two by randomly choosing $n_{\mathrm{layer},\mathrm{c}}\in\{1,2,3,4\}$ dense layers for each neural network, and independently choosing $n_{l,\mathrm{c}}\in\{2,3,4\}$ neurons for each dense layer at random, with $l=1,\dots,n_{\mathrm{layer},\mathrm{c}}$. Note that we actively decide to keep both numbers small in order to keep the models simple and the number of trainable parameters of the ensuing classical neural networks comparable to those of the quantum neural networks whose randomized architectures are introduced in Sec. III.1.2. While the number of trainable parameters is only one factor among many defining the (expressive) power of machine learning models, with other factors being, for instance, their non-linearities, it is nevertheless a vital factor and its equivalence a prerequisite for comparability between different models — i.e., between classical and quantum architectures in our case. We thereby diminish the possibility that potentially better results for either classical or quantum neural networks are due to the fact that one has more trainable parameters available than the other. For the activation functions, we restrict our study to architectures where we use ReLU [36], defined as $\mathrm{ReLU}(x)=\max(0,x)$, between every layer except for the final layer, which consists only of a single neuron for classification that is connected to its previous layer via the sigmoid activation function, defined as $\mathrm{\sigma}(x)=(1+e^{-x})^{-1}$.
@@ III.1.2 Quantum neural networks
In the following, we introduce how we generate random quantum neural network architectures, which, in practice, are given by parametrized quantum circuits (PQCs). To this end, note that their hyperparameters are — at least partially — very different to those of classical neural networks. As mentioned earlier,
@@
In order to adapt the above scheme to our approach of generating random quantum neural network models, there are various hyperparameters that can be varied. First, we allow for a random number $n_{\mathrm{q}}$ of qubits. In detail, to generate a random quantum neural network able to process $d$-dimensional input vectors, we choose $n_{\mathrm{q}}\in\{2,4,8,\dots,d\}$ randomly, exploiting that $d$ is always a power of two. Second, and intimately connected to the choice of $n_{\mathrm{q}}$, is the choice of the quantum feature map $\mathsf{S}_{\mathrm{enc}}(\boldsymbol{x}_{i})$, which, inspired by Refs. [24, 39, 40], may become trainable as well in our case, i.e., $\mathsf{S}_{\mathrm{enc}}(\boldsymbol{x}_{i},\boldsymbol{\phi})$ with $\boldsymbol{\phi}\in\mathbb{R}^{M_{\mathrm{enc}}}$ a vector of $M_{\mathrm{enc}}$ trainable parameters. For simplicity, we first discuss the case $n_{\mathrm{q}}=d$, since the case $n_{\mathrm{q}}<d$ is more contrived. For $n_{\mathrm{q}}=d$, the quantum feature map is chosen such that
@@
where $\boldsymbol{\phi}_{l}$ and $\boldsymbol{\theta}_{l}$ are trainable parameter that are independent for each layer $l$ despite $\mathsf{S}_{\mathrm{enc}}(\boldsymbol{x}_{i},\boldsymbol{\phi}_{l})$ and $\mathsf{U}(\boldsymbol{\theta}_{l})$ having the same overall structure in each layer. As a final choice of randomizing the PQC, the additional unitary $\mathsf{U}_{\mathrm{had}}$ randomly applies either a Hadamard gate or the identity to all qubits. This is a means to rotate the initial state |$\ket{\Psi_{\mathrm{init}}}=\ket{0}^{\otimes n_{\mathrm{q}}}$ into a more accessible basis prior to executing the remainder of the PQC.

### v2

@@ Abstract
Image classification tasks are among the most prominent examples that can be reliably solved by classical machine learning models. In this study, we compare the performance of randomized classical and quantum neural networks as well as classical and quantum-classical hybrid convolutional neural networks for the task of binary image classification. We employ two distinct methodologies: (1) using randomized (quantum) neural networks on dimensionality-reduced data, and (2) applying (hybrid) convolutional neural networks to full image data. We evaluate these approaches on three data sets of increasing complexity: (i) an artificial hypercube dataset, (ii) MNIST handwritten digits (0’s and 1’s), and (iii) real-world industrial images from laser cutting machines. We analyze correlations between classification accuracy and quantum model hyperparameters, including the number of trainable parameters, feature encoding methods, circuit layers, entangling gate type and structure, gate entangling power, and measurement operators. This provides a systematic hyperparameter optimization study rarely conducted in quantum machine learning literature. For random quantum neural networks, we compare their performance with three literature models: Chebyshev [1], Hubregtsen [2] and HighDim [3]. Quantitatively, classical and quantum/hybrid models achieved statistically equivalent classification accuracies across most datasets, with no single approach demonstrating consistent superiority. We observe that quantum models show lower variance with respect to initial training parameters, suggesting better training stability. Among the hyperparameters analyzed, only the number of trainable parameters showed a consistent positive correlation with the model performance. Around $94\%$ of the best-performing quantum neural networks had entangling gates, although for hybrid convolutional neural networks, models without entanglement performed equally well but took longer to converge. Cross-dataset performance analysis revealed limited transferability of quantum models between different classification tasks. The inconsistent importance of entanglement across our experiments highlights the current limited theoretical understanding of how quantum models actually function and what determines their performance. Our study provides an industry perspective on quantum machine learning for practical image classification tasks, highlighting both current limitations and potential avenues for further research in quantum circuit design, entanglement utilization, and model transferability across varied applications.
@@ I Introduction
Quantum machine learning [4, 5] is one of the various subfields of quantum computing promising advantages over its classical counterpart, for instance due to proposals regarding more efficient data encoding [6, 7], improved model expressivity [8, 9, 10], prediction accuracy [11, 12, 13] or by saving computational resources [14, 15]. These theoretical proposals have been backed up by claims of experimental quantum advantage in learning tasks [16, 17]. Despite undeniable progress in the field of quantum machine learning — both on the algorithmic as well as the hardware side — doubts about the field’s focus on quantum advantage have been previously raised [18]. Ultimately, this culminated in a recent work questioning the significance of quantum-machine-learning benchmark studies [19], given their predominant use of overly simple data sets paired with insufficiently-performing quantum hardware. Other works [20, 21, 22] even suggest that much of what was commonly believed to be classically hard or intractable to simulate might be efficiently simulable on classical hardware after all. While this may require the field to reevaluate what “quantum” means for quantum machine learning and how a future quantum computer may use “quantumness” advantageously for learning tasks, both Refs. [19, 20] independently call for more rigorous benchmarks employing less trivial data sets in order to assess the potential of quantum machine learning. In response to this critical need, our work directly addresses this challenge by employing not only standard benchmark datasets but crucially incorporating real-world industrial data that represents practical machine learning applications beyond academic test cases.
@@
In this work, we benchmark the performance of classical and quantum machine learning models in classifying three data sets of increasing data complexity. The data sets include (i) an artificial data set based on linearly splitting a (hyper-)cube in two, (ii) MNIST’s handwritten digits [23] as an easy but well-studied data set and (iii) images from laser cutting machines as a challenging data set from the industry. Note that we use the first two data sets merely as a means to benchmark and validate our employed methodologies, since — as mentioned in Refs. [19, 21] — the potential insights about the power and usefulness of quantum machine learning that can be drawn from simple data sets are very limited. However, we still believe them to be useful for validating the results we obtain for the practically-relevant industrial data set on which we put the main focus when drawing conclusions on quantum machine learning from an industry perspective. Moreover, since we are also interested in testing the cross-data-set performance, assessed by the classification accuracy, of some quantum models, we deem it beneficial to have results for various data sets with different levels of complexity available.
@@
In terms of methodologies, we employ both classical and quantum neural networks [8] as well as classical and quantum-classical hybrid convolutional neural networks [24] to classify any of the mentioned data sets. The dual-methodology approach used allows us to systematically analyze not just the relative performance of classical versus quantum models, but also to extract insights on which quantum architecture aspects contribute most to performance, how hyperparameters correlate with classification accuracy, and whether successful quantum models on one dataset transfer well to others — providing a comprehensive framework for quantum advantage assessment in realistic settings. Our decision for parametrized quantum circuits is motivated by their high expressivity [8] which is amplified by techniques like data re-uploading [25, 26] and trainable data embeddings [27] while at the same time keeping quantum hardware requirements moderate. This is in contrast to some early suggestions for quantum machine learning models based on amplitude encoding [7, 6, 28], which, despite being qubit-efficient, were later deemed too resource-demanding to run on any near-term device due to their excessive need of quantum gates for accurate state preparation [9]. In contrast to these advantages, parametrized quantum circuits suffer from barren plateaus [29], which prevents their efficient training at scale, although much work has been put into mitigating or avoiding them, see e.g. Refs. [30, 31, 32, 33] to name but a few. Some quantum models, such as quantum convolutional neural networks [34], provably avoid barren plateaus [35] but are classically efficiently simulable [21]. Unfortunately, the latter are also rather resource-demanding to qualify for near-term applicability and practical simulability within this study, which is why we resort to quantum-classical hybrid convolutional neural networks [24] instead. Conceptually, parametrized quantum circuits as well as quantum-classical hybrid convolutional neural networks belong to the realm of variational quantum algorithms [36], where the execution of quantum circuits is carried out on actual quantum hardware, while the calculation of iteratively-updated circuit parameters is performed on classical computers. Due to these choices, our study reflects a current perspective on what quantum machine learning methods, suitable for near-term quantum devices, offer for classifying industry-level image data.
@@ II Data sets
Figure 1: Overview of data sets with $N=500$ and dimension $d=2$ in panels (a)-(e), and the corresponding learned functions by classical and quantum neural networks in panels (f)-(j) and panels (k)-(o), respectively. The small neural networks in between panels (a)-(e) and (f)-(j) show the best performing architectures (from among the set of randomly generated neural networks). The small quantum neural networks in between panels (a)-(e) and (k)-(o) show the respective best performing quantum architectures (from among the set of randomly generated quantum neural networks). Data points represented by circles (stars) in panels (a)-(o) correspond to the training (validation) data set. In addition, panels (p) and (q) show the architecture for the convolutional autoencoder (CAE) used to reduce the data dimensionality for the images of the second and third data set, respectively, as well as a representative illustration of the original and reconstructed images for various latent dimensions $d$.
@@ II.1 Data set 1 — linear split
The first data set is an artificial one and the only one not based on images. It is chosen with the premise of simplicity in mind and such that the ideal function, which classifies the data set perfectly, is known in advance. To this end, we choose to consider a $d$-dimensional hypercube, defined by the set $[0,1]^{d}\subseteq\mathbb{R}^{d}$, and linearly split it by the plane orthogonal to the normal vector $\boldsymbol{n}=(1,1,\dots,1)^{\top}/\sqrt{d}\in\mathbb{R}^{d}$ which passes through the hypercube’s center $(1,1,\dots,1)^{\top}/2\in\mathbb{R}^{d}$. In order to generate training and validation data sets, we first place $N$ data points randomly within the hypercube such that exactly $N/2$ points are located within each of the two sectors separated by the hyperplane. These two sectors define the two labels that classify each data point. Afterwards, we split the $N$ data points into training (validation) data sets consisting of $N_{\mathrm{train}}=0.8N$ ($N_{\mathrm{val}}=0.2N$) of the total data points. We perform this splitting so that the number of data points with each label within each subset is identical. In order to study the effect of the hypercube’s dimension $d$ and the data set size $N$, we generate versions of the hypercube data set with $d\in\{2,4,8,16,32,64\}$ and $N\in\{200,500,1000,2000\}$. An example with $d=2$ and $N=500$ is shown in Fig. 1 (c).
This hypercube data set has recently also been studied in a related work exploring the performance of quantum machine learning methods [19]. In this work, the hypercube data turned out to be rather difficult to classify by means of quantum machine learning methods, despite being a trivial problem for classical machine learning. Hence, we use it in our study as an ideal means to validate and benchmark our methodologies, introduced in Sec. III, and to check whether it can reproduce those results qualitatively.
@@ II.2 Data set 2 — handwritten digits
As a second data set, we take the well-studied handwritten digits of the MNIST database [23] as another means to test our methodologies. However, in order to keep the classification problem binary, we restrict our study to images of 00’s and $1$’s. Similar to the hypercube’s data sets described in Sec. II.1, we generate different versions of the MNIST data set with $N\in\{200,500,1000,2000\}$ total images taken randomly from the $60000$ available images. We again ensure that each generated data set has an equal number of 00’s and $1$’s and that this ratio is maintained once the total data set is split into training and validation sets with $N_{\mathrm{train}}=0.8N$ and $N_{\mathrm{val}}=0.2N$ data points, respectively.
@@ II.3 Data set 3 — industrial images
In contrast to Fig. 2, the slat images we use in the following are already preprocessed and cut such as to focus on a single slat socket which may or may not contain a slat. The classification problem is thus binary with two difficulties arising due to varying camera angles from which the slat sockets are being photographed as well as the different condition of the slats, cf. Fig. 2. To ease data generation, TRUMPF has developed a model that generates synthetic slat socket images given the CAD model of the entire laser cutting machine and textures derived from real-world images using the Blender rendering engine. This way, manual labeling effort of real-world images is unnecessary and labeling errors are ruled out. Moreover, it allows to generate data sets with large image numbers. Identically to the other data sets, we consider different versions of this industry data set with $N\in\{200,500,1000,2000\}$ total images, which are afterwards split into training and validation data sets with $N_{\mathrm{train}}=0.8N$ and $N_{\mathrm{val}}=0.2N$ data points, respectively. For these two sets, we ensure again the same class distribution as the original data set, as well as similar number of images of slat sockets photographed with different camera angles and conditions. All images are gray-scale and normalized to $200\times 80$ pixels independent of the photo angle.
@@ III.1.1 Classical neural networks
In order to generate random classical neural network architectures that allow processing of data vectors $\boldsymbol{x}_{i}\in\mathbb{R}^{d}$, there are various hyperparameters that can be randomly determined. Among the most prevalent ones are the number of dense, i.e., fully-connected, layers and the number of neurons per layer as well as the employed activation functions. We decide to randomize the first two by randomly choosing $n_{\mathrm{layer},\mathrm{c}}\in\{1,2,3,4\}$ dense layers for each neural network, and independently choosing $n_{l,\mathrm{c}}\in\{2,3,4\}$ neurons for each dense layer at random, with $l=1,\dots,n_{\mathrm{layer},\mathrm{c}}$. Note that we actively decide to keep both numbers small in order to keep the models simple and the number of trainable parameters of the ensuing classical neural networks comparable to those of the quantum neural networks, whose randomized architectures are introduced in Sec. III.1.2. While the number of trainable parameters is only one factor among many defining the (expressive) power of machine learning models, with other factors being, for instance, their non-linearities, it is nevertheless a vital factor and its equivalence a prerequisite for comparability between different models — i.e., between classical and quantum architectures in our case. We thereby diminish the possibility that potentially better results for either classical or quantum neural networks are due to the fact that one has more trainable parameters available than the other. For the activation functions, we restrict our study to architectures where we use ReLU [39], defined as $\mathrm{ReLU}(x)=\max(0,x)$, between every layer except for the final layer, which consists only of a single neuron for classification that is connected to its previous layer via the sigmoid activation function, defined as $\mathrm{\sigma}(x)=(1+e^{-x})^{-1}$.
@@ III.1.2 Quantum neural networks
In the following, we introduce how we generate random quantum neural network architectures, which, in practice, are given by parametrized quantum circuits (PQCs). To this end, note that their hyperparameters are — at least partially — very different to those of classical neural networks.
@@
In order to adapt the above scheme to our approach of generating random quantum neural network models, there are various hyperparameters that can be varied. First, we allow for a random number $n_{\mathrm{q}}$ of qubits. In detail, to generate a random quantum neural network able to process $d$-dimensional input vectors, we choose $n_{\mathrm{q}}\in\{2,4,8,\dots,d\}$ randomly, exploiting that $d$ is always a power of two. Second, and intimately connected to the choice of $n_{\mathrm{q}}$, is the choice of the quantum feature map $\mathsf{S}_{\mathrm{enc}}(\boldsymbol{x}_{i})$, which, inspired by Refs. [27, 2, 42, 43], may become trainable as well in our case, i.e., $\mathsf{S}_{\mathrm{enc}}(\boldsymbol{x}_{i},\boldsymbol{\phi})$ with $\boldsymbol{\phi}\in\mathbb{R}^{M_{\mathrm{enc}}}$ a vector of $M_{\mathrm{enc}}$ trainable parameters. For simplicity, we first discuss the case $n_{\mathrm{q}}=d$, since the case $n_{\mathrm{q}}<d$ is more contrived. For $n_{\mathrm{q}}=d$, the quantum feature map is chosen such that
@@
where $\boldsymbol{\phi}_{l}$ and $\boldsymbol{\theta}_{l}$ are trainable parameters that are independent for each layer $l$ despite $\mathsf{S}_{\mathrm{enc}}(\boldsymbol{x}_{i},\boldsymbol{\phi}_{l})$ and $\mathsf{U}(\boldsymbol{\theta}_{l})$ having the same overall structure in each layer. As a final choice of randomizing the PQC, the additional unitary $\mathsf{U}_{\mathrm{had}}$ randomly applies either a Hadamard gate or the identity to all qubits. This is a means to rotate the initial state |$\ket{\Psi_{\mathrm{init}}}=\ket{0}^{\otimes n_{\mathrm{q}}}$ into a more accessible basis prior to executing the remainder of the PQC.